OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and perpendicular to y = 1 fifthx - 10. y = 5x + 10 y = 5x y = -5x - 10 y = -5x

OpenStudy (anonymous):

the product of slopes of perpendicular lines =-1 y=(1/5)x-10 so the slope of perpendicular line : -5 y-y1=m(x-x1) y+5=-5(x-1) y+5=-5x+5 y=-5x+5-5 y=-5x